NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 09:31:28 on February 27, 2009
Entry number 263668
This entry is a followup to: 263544
Followups:
keyword=Experiment Summary for Feb 26, Owl - Feb 26, Swing
Summary for Feb 26 owl - Feb 26 swing
Smooth data taking at fourth pass, then pass change to fifth.
Beam in Use
shift |
ABU |
BANU |
BNA |
ACC |
2009/02/26 owl | 6.53 |
0.55 | 0.9 | 0 |
2009/02/26 day | 0.34 |
1.31 | 5.66 | 0.69 |
2009/02/26 swing | 4.18 |
1.99 | 1.82 | 0 |
total | 11.1 |
3.8 | 8.4 | 0.7 |
Main events/issues:
- February 26th
- Completed production on kinematic 4J: Plhrs = 1.51
GeV/c, 45o.
- 02:00 Change to kinematic 4C: Plhrs to 0.8 GeV/c,
45o.
- 08:22 Beam studies start early due to North LINAC vacuum problem.
- Accumulated charge kinematic 4J: 1.14 C out of 1.08 C.
- Accumulated charge kinematic 4C: 0.25 C out of 0.38 C.
- 8:30 Controlled access:
- Spare reference cell mounted.
- HRS magnet work: Right Q3 now operational and some work on left Q3.
- BigBite gas Cerenkov gains checked.
- Target spin rotated to 270o.
- Change to kinematic M: Plhrs to 1.7 GeV/c, 45o.
- 10:15 Wien angle change to 30o and spot move on
photocathode.
- 16:20 Beam tuning for 5th pass begins.
- 18:00 Beam tuning complete.
- 18:00-18:35 Controlled access for BigBite Cerenkov missing TDC
channel.
- 19:00-20:30 N2 reference cell runs.
- 19:35 Change to kinematic K: Plhrs to 1.51 GeV/c,
45o.
- 19:49 Begin production for kinematic K.
Plans for the next couple of days:
- February 27th to March 2nd: Production on
kinematics K, I, G, E and C.
- March 2nd:
- 6:30 Controlled Access for left HRS polarity change to +ve.
- 9:00-17:00 left HRS positron data.
- 17:00 Controlled access for left HRS polarity change to -ve.
- 19:00 Moller measurement.
- 23:00 Resume polarized production at 5th pass.
A copy of this log entry has been emailed to: brads,jpchen,jiang,meziani,choi